OpenStudy (anonymous):

A program on television ran 30 minutes. The program ran 7 minutes longer than 3 times the length of the commercials. Let the length of the commercials be represented by the variable x. Which equation represents this problem? A. 3x + 7 = 30 B. x + 7 = 3(30) C. 3x = 30 + 7 D. 7x + 3 = 30

OpenStudy (♪chibiterasu):

30 is the main product, so we first have: = 30 3 times the length of the commercials means 3(x), x being whatever the commercials last in terms of time. 3x = 30 << not complete 7 minutes longer means we are going to add 7 on to the soooo: 3x + 7 = 30 Success :D To find the length of the commercials just solve: 3x/3 + 7 = 30/3 x + 7 = 10 x + 7 - 7 = 10 - 7 x = 3 The commercials last 3 minutes :3

OpenStudy (anonymous):

so its b?